OpenStudy (anonymous):

The quadrilateral Formed by the Straight line \[x \pm y \pm 1 =0 \]

OpenStudy (shubhamsrg):

x+y +1 = 0 x+y -1 = 0 x-y +1 =0 x-y -1 = 0 where are you stuck ?

OpenStudy (anonymous):

No..Idea....!

OpenStudy (shubhamsrg):

you should be able to see 2 pairs of // and perpendicular lines .. why dont you try and plot em ?

OpenStudy (anonymous):

It may be Square or Rectangle..)

OpenStudy (shubhamsrg):

yep..

OpenStudy (shubhamsrg):

if you see properly those eqns in intercept form, you should be able to identify if its a rectangle or sq.